Costumeers arrive at a checkout counter in a department store according to a Poisson distribution

at an average of seven per hour. During a given hour, what are the probabilities that

(a) no more than three customers arrive?

(b) at least two customers arrive?

(c) exactly four customers arrive?



1
Expert's answer
2021-05-07T09:23:59-0400

Let "X=" the number of costumers arrive at a checkout counter in a department store: "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=7."

(a)

"P(X\leq3)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)=\dfrac{e^{-7}\cdot7^0}{0!}+\dfrac{e^{-7}\cdot7^1}{1!}+\dfrac{e^{-7}\cdot7^2}{2!}"

"+\dfrac{e^{-7}\cdot7^3}{3!}=e^{-7}(1+7+\dfrac{49}{2}+\dfrac{343}{6})"

"\approx0.081765"

(b)


"P(X\geq2)=1-(P(X=0)+P(X=1))"

"=1-\dfrac{e^{-7}\cdot7^0}{0!}-\dfrac{e^{-7}\cdot7^1}{1!}=1-8\cdot e^{-7}"

"\approx0.992705"

(c)


"P(X=4)=\dfrac{e^{-7}\cdot7^4}{4!}\approx0.091226"